OpenStudy (nikkiv73):

Find the 20th term of the following sequence. -6, -4, -2, 0, ... 32 44 46

OpenStudy (nikkiv73):

I feel like it should be 24 but that's not an option. What am I missing?

OpenStudy (sylph):

You're counting by twos. -6, -4, -2, 0, 2, 4, 6, 8....

OpenStudy (sylph):

What would be the twentieth number in that line?

OpenStudy (nikkiv73):

So then wouldnt the 20th term be 16? Which also isn't' an option

OpenStudy (sylph):

-6, -4, -2, 0, 2, 4, 6, 8, 10, 12, 14, 16.... Is that 20 numbers?

OpenStudy (nikkiv73):

ahhh 32!! Thank you - i should have written it down for a visual

OpenStudy (nikkiv73):

What is the formula to figure this out?

OpenStudy (sylph):

Whoops. \[20\times2-8\]

OpenStudy (sylph):

>-<

OpenStudy (nikkiv73):

why minus 8?

OpenStudy (sylph):

Because \[-6+-4+-2+0>1\]

OpenStudy (nikkiv73):

OK thanks!

OpenStudy (sylph):

Each one has a difference of two, meaning they are four irrelevant twos. Your 20th digit will not include them. No problem!

OpenStudy (plainntall):

It looks like you are doubling the term number and then subtracting by 6 with n starting at 0. 2n-6 (2(0)-6)=-6 for n=0 first term (2(1)-6)= -4 for n=1 second term (2(2)-6) = -2 for n=2 third term (2(3)-6)= 0 for n-3 4th term you can keep on adding 2 to the previous term until you have the 20th term (which would be n=19 since we started at n=0) or use the equation we intuited 2n-6 = (2(19)-6) = 38-6 = 32

OpenStudy (nikkiv73):

Thank you!